### v2

# Can We Catch the Elephant? A Survey of the Evolvement of Hallucination Evaluation on Natural Language Generation

Abstract
Hallucination in Natural Language Generation (NLG) is like the elephant in the room, obvious but often overlooked until recent achievements significantly improved the fluency and grammaticality of generated text. As the capabilities of text generation models have improved, researchers have begun to pay more attention to the phenomenon of hallucination. Despite significant progress in this field in recent years, the evaluation system for hallucination is complex and diverse, lacking clear organization. We are the first to comprehensively survey how various evaluation methods have evolved with the development of text generation models from three dimensions, including hallucinated fact granularity, evaluator design principles, and assessment facets. This survey aims to help researchers identify current limitations in hallucination evaluation and highlight future research directions.
[…]
Can We Catch the Elephant? A Survey of the Evolvement of Hallucination Evaluation on Natural Language Generation
Siya Qi, Yulan He, Zheng Yuan King’s College London, London, UK {siya.qi, yulan.he, zheng.yuan}@kcl.ac.uk
1 Introduction
Hallucination in the context of Natural Language Generation (NLG) typically refers to situations where the generated text is inconsistent with, or unsupported by, the input source or external knowledge. Like an elephant in the room, this problem has existed since the beginning of the development of NLG. As text generation models continue to evolve, technologies like Large Language Models (LLMs) have achieved grammatical correctness and fluency nearly indistinguishable from human writing. Consequently, hallucination has gradually surfaced and attracted increased attention. The evaluation of hallucination matters since it motivates the development of LLMs to be more reliable and safe. In this paper, we make a comprehensive survey on the evolution of hallucination evaluation methods through multiple perspectives as in Figure 1, which gives a broad view of current trends and future directions.
The definition of hallucination varies a lot from work to work over time. In the initial stages, hallucination referred to the occurrence of semantically incorrect but grammatically viable content (Lee et al., 2018). This setup was typically rooted in traditional tasks such as Summarization and Neural Machine Translation (NMT), which often involve definite source information. The breakthrough came with the advent of LLMs like ChatGPT (OpenAI, 2022), crossing the barriers of conversational systems. These models are capable of processing various text inputs and providing responses. In this case, users have more expectations for the responses, especially the consistency of world facts under open answers. Many previous tasks can also be accomplished within a conversation by prompting LLMs with correct instructions (Ouyang et al., 2022). In the responses of LLMs, hallucination occurs now and then, with a mix of faithful and factual problems.
Figure 1: The distribution of evaluators on various aspects, including the timeline, hallucinated fact granularity, method design principles, and assessment facets.
Faithfulness and factuality are two concepts that are closely related when describing hallucinations in generated text and can be prone to confusion in some circumstances. In this paper, we add prefixes to both of them for better understanding by introducing Source Faithfulness (SF) and World Factuality (WF). [^1] In this paper, we also write SF as faithfulness/consistency, and WF as factuality. SF measures the degree to which the generated output accurately reflects and is consistent with the source input. It ensures that the output stays true to the information provided in the source. SF has a limited scope, as there are specific sources that can be used to substantiate and verify the generated text. WF, on the other hand, assesses whether the generated output aligns with general world knowledge and facts. WF is a more expansive and challenging problem as it goes beyond the specific source and considers the broader context of common sense and established knowledge, which is more difficult to collect and encode comprehensively. Recent studies have recognized the critical importance of addressing and improving the SF and WF of generated text.
Assessing from SF or WF aspects means the evaluators refer to different source information, which is closely tied to specific tasks. For example, in NMT, generated translations detached from the source text are considered factually incorrect (Dale et al., 2023). In summarization, summaries usually should be faithful to the source document, but some also argue that certain hallucinations can align with external facts (Dong et al., 2022). In dialogue generation, it is important to balance hallucination and user experience (Santhanam et al., 2021). In LLM downstream tasks, the issue is often related to the prompts, which are designed for different scenarios and to retrieve the knowledge that exists in LLMs. Apart from these, LLMs face unique difficulties, such as updating world information and handling false-premise questions (Vu et al., 2023; Kasai et al., 2024).
Hallucinations are prevalent across various tasks, as the problem has not been completely defined, yet it intuitively pertains to binary correctness. While humans can readily perceive some hallucinations, models often struggle to do so. Moreover, in the current development of LLMs regarding reasoning, planning, task adaptability, etc., hallucination evaluation is an issue that requires careful consideration. Previous works have some introductions on benchmarks for LLM hallucinations (Huang et al., 2023; Zhang et al., 2023c; Ji et al., 2023; Huang et al., 2021), but they do not include evaluators before LLMs and more recent research, thus resulting in a lack of comparison between developments before and after. Also, they do not specifically focus on evaluators, therefore lacking in-depth analysis of details, as well as their strengths and weaknesses.
Our contributions lie in three folds. Firstly, this paper organizes automatic evaluation methods for hallucination [^2] We also give a brief introduction of related fact-checking and human evaluation works in Appendix A and B. in the evolution timeline, including traditional approaches as well as novel methods employing or assessing LLMs, and conducts analytical comparisons. Secondly, this paper reviews various evaluation methods from three perspectives: hallucinated fact granularity, evaluator design principles, and assessment facets, highlighting the intrinsic connections and development trajectories among these methods. Finally, we summarize the current state of research on hallucination evaluation, outlining existing issues and potential future directions.
2 Fact Definition
To evaluate hallucination in text generation, we first need to define what constitutes a fact. Previous works have evaluated hallucination by considering different levels of fact granularities and categorizing fact errors into various types.
2.1 Fact Granularity
The definition of fact granularity varies, lacking an agreed-upon labeling scheme (Honovich et al., 2022), and many elements in generated text are difficult to define clearly. Previous works categorize fact granularity into the following types.
Figure 2: Source Faithful Error and World Factual Error examples.
Token/Word: At the token/word level, entities and nouns are the most frequently used granularities, primarily due to their intuitively informative nature and relative ease of acquisition and comparison. For example, some summarization models tend to forget (Maynez et al., 2020; Nan et al., 2021) or make up entities that never appeared in the source document.
Span: Spans require more annotation than individual tokens or words. Mostly, they are entity-centric, relation-centric, or event-centric, closely related to dependency parsing (Goyal and Durrett, 2020, 2021). In evaluators, spans can be used to generate questions (Durmus et al., 2020; Honovich et al., 2021) or check entailment (Kamoi et al., 2023).
Sentence: A higher level of granularity is the sentence, which is coarsely classified as containing hallucination or not (Wang et al., 2020; Pagnoni et al., 2021). Sentence-level facts provide an overall perception of the factual content within the sentence, allowing for a more holistic assessment.
Passage: Passage-level facts contribute to an overall assessment of the generated text (Fabbri et al., 2021), encompassing discourse fact connections (Zhang et al., 2023b) and an estimation of passage themes. This level of granularity is crucial for evaluating longer texts.
Figure 1 illustrates that evaluation methods before the advent of LLMs mostly relied on text segments, whereas many methods developed after LLMs frequently conduct direct assessments of the entire input text.
2.2 Fact Error Type
The taxonomies of fact errors vary depending on the scenarios. Some categorize them by the information source as intrinsic and extrinsic hallucinations (Maynez et al., 2020; Goyal and Durrett, 2021; Cao and Wang, 2021). For faithful error types, Goyal and Durrett (2021) designs entity-, event-, and noun phrase-related errors, while Pagnoni et al. (2021) categorize errors on a more fine-grained scale, including PredE, EntE, CircE, CorefE, LinkE, OutE and GramE. We propose categorizing errors into SF Error and WF Error, which can cover a wider range of NLG task, including both traditional ones and LLM scenarios. SF and WF errors can be assessed independently, as illustrated in Figure 2.
3 Evaluation Methods Before LLMs
This section discusses automatic evaluation methods for hallucinations in NLG before the era of LLMs. These methods differ from those that focus on text similarity (Lin and Hovy, 2003) or grammatical correctness (Bryant et al., 2017), as their primary focus lies on the SF aspect of text generation. Previous works rely on references to train models, but these references can contain factual errors (Maynez et al., 2020), hindering the model’s ability to learn faithfulness. Due to the costly and error-prone nature of references, many works study reference-free evaluation methods that only require the input of source documents.
A significant portion of these evaluation methods is dedicated to assessing the summarization task, which has a fixed source document as an information reference, making faithfulness relatively straightforward to evaluate. Summarization is commonly employed in fields such as news, medicine, and law, where reliability is of particular importance. In addition to summarization, this section covers tasks such as text simplification, NMT, and dialogue generation. Some evaluation methods concentrate on a single task, while others can be extended to similar ones. In the following subsections, we will introduce the approaches by elaborating on evaluators and benchmarks.
3.1 Reference-based Evaluators
Reference-based evaluators compare generated text with ground-truth references to assess factual consistency. For summarization, early approaches start with statistical methods. $\mathbf{Fact_{acc}}$ (Goodrich et al., 2019) uses information extraction models to measure factual overlap based on pre-defined schemas. FactCC and FactCCX Kryscinski et al. (2020) use Natural Language Inference (NLI) models for fact classification and span highlighting. Besides utilising entity-level facts, DAE (Goyal and Durrett, 2020) focuses on dependency-level factual error using entailment models.
Text simplification aims to “translate” texts to a more readable version without altering their meaning. Maskeval (Liu et al., 2022) evaluates both summarization and simplification by combining masked LM weights with Exact Matching (EM) scores, exploiting the similarity between the two tasks.
Hallucinations in NMT have received attention as well (Lee et al., 2018; Raunak et al., 2021). Hallucination in NMT is rare, like finding a needle in a haystack. It is important to identify the hallucinations with the errors.  Guerreiro et al. (2023) propose a glass-box uncertainty-based detection method, while  Dale et al. (2023) measure hallucination by computing the percentage of source contribution to the generated translation.
These approaches provide promising solutions, as they explore existing datasets and deliver feasible solutions. For long-input tasks such as summarization and simplification, they can use label information, with manual correction making it more accurate. However, a notable drawback is the requirement for annotation, which is costly and labor-intensive, and may also need cross-checking. Additionally, distinguishing between normal translation errors and hallucinations in NMT presents significant challenges.
3.2 Reference-free Evaluators
Reference-free evaluation methods assess the quality of generated text without relying on direct ground truth. For abstractive summarization, the Question Generation and Question Answer (QG-QA) pipeline is a common approach. FEQA (Durmus et al., 2020), QAGS (Wang et al., 2020), and QuestEval (Scialom et al., 2021) follow a similar framework: extracting question-answer pairs from the summary and then using off-the-shelf QA models to check the consistency of answers extracted from the document. QuestEval also evaluates four dimensions beyond faithfulness and uses a weighter to identify important questions. QAFactEval (Fabbri et al., 2022) analyzes and combines the best components of the QA-based metric pipeline for improved performance, including answer selection, question generation, question answering, answer overlap evaluation, and question filtering. MQAG (Manakul et al., 2023a) computes the statistical distance of answers over automatically generated multiple-choice questions, allowing for a more abstract understanding of information. Counterfactual estimation, such as CoCo (Xie et al., 2021), measures the causal effects from source documents to generated summaries. It can also be generalized to multi-level granularity and other tasks. Other than QA-QG pipeline, FactGraph (Ribeiro et al., 2022) compares the generated summary with the input document through semantic graphs, thereby enhancing discourse comprehensibility. FactKB (Feng et al., 2023) introduces external triplet knowledge to augment the training data, and assess the consistency of news summarization. ExtEval (Zhang et al., 2023b) reveals that extractive summarization does not guarantee faithfulness, identifying five broad categories of unfaithfulness issues that the entailment model cannot fully address.
For Dialog Generation, $\mathbf{Q^{2}}$ (Honovich et al., 2021) is a evaluation method for knowledge-grounded generative models. It incorporates QG, QA, and NLI models to measure the model’s output to reflect its consistency with its grounding source knowledge, and measures the answer span similarity using an NLI model. FactPush (Steen et al., 2023) provides a simple data augmentation method by adding phrases to sentences, which can improve the NLI model performance.
For multiple tasks, AlignScore (Zha et al., 2023) can give a unified function for 7 tasks consistency evaluation. It breaks the context into chunks and the claim into sentences to calculate the alignment score. WeCheck (Wu et al., 2023) provides a weak supervision learning paradigm based on previous metrics and uses a mixture NLI datasets to warm up and do a noise-aware fine-tuning to get the target metric model.
Compared to reference-based methods, these approaches have greater generality. It is noticeable that QG-QA models take a large portion. Because, without references, it is challenging for NLI models to get accurate extraction of evidence from the source text. Essentially, the QG-QA framework is a flexible approach that extracts and compares information from generated output and source input. However, the QG and QA steps may introduce considerable noise, which can interfere with the results. When comparing, some methods also incorporate entailment models for judgment. It is showed that entailment and QA-based metrics offer complementary signals that the combination can achieve better performance (Fabbri et al., 2022).
3.3 Datasets and Benchmarks
Alongside evaluation models, datasets and benchmarks have been proposed to measure hallucination across tasks and assess evaluator efficacy.
For summarization,  Maynez et al. (2020) find severe extrinsic hallucination in the reference of XSum dataset, and label XSumFaith with hallucination type (intrinsic or extrinsic) and hallucination span position. CoGenSumm (Falke et al., 2019) is collected from the validation set of CNN/DM and is used to rerank generated summaries. From both CNN/DM and XSum, QAGS (Wang et al., 2020) annotates each sentence with a binary choice based on whether the sentence is faithful to the article. However, the binary factuality of texts can be difficult to determine. FRANK (Pagnoni et al., 2021) collect annotation based on a more fine-grained defined typology of factual errors. Polytope (Huang et al., 2020) provides both SF and WF annotations to measure both extractive and abstractive summarization. Falsesum (Utama et al., 2022) mimics intrinsic errors and extrinsic errors to augment summary datasets for NLI models.  Devaraj et al. (2022) label 3 types of factual errors for data collected from Newsela (Xu et al., 2015) and Wikilarge (Zhang and Lapata, 2017) for text simplification. In dialogue generation, Dialogue NLI (Welleck et al., 2018) provides three-type labels for NLI models of the entailment of sentence pairs.
Given the large amount of evaluation metrics proposed, some benchmarks are designed to reevaluate these metrics. SummEval (Fabbri et al., 2021), SummaC (Laban et al., 2022), Dialsummeval (Gao and Wan, 2022), and AGGREFACT (Tang et al., 2023) focus on summarization hallucination evaluation or detection. Wizard of Wikipedia (Dinan et al., 2018), CI-ToD (Qin et al., 2021), BEGIN (Dziri et al., 2022b), FaithDial (Dziri et al., 2022a) and TopicalChat (Gopalakrishnan et al., 2023) can help measure consistency evaluators on dialogue generation. Beyond a single task, TRUE (Honovich et al., 2022) and BEAMetrics (Scialom and Hill, 2021) can evaluate metrics across a series of text generation tasks.
4 Evaluation Methods After LLMs
With the rapid development of LLMs and their human-like performance, LLMs possess instruction-tuned LLMs, like ChatGPT (OpenAI, 2022) or GPT-4 (Achiam et al., 2023), which offer an advantage that can take instruction in natural language as human experts when evaluating. For SF and WF in LLM scenarios, they are mainly presented in the form of context consistency or factual imagination. Some works prompt LLMs as direct evaluators, while other works design a pipeline with multiple steps to do the assessment. In this section, we will elaborate on both aspects.
4.1 Direct Evaluation
Recent LLMs are powerful tools that can be prompted by instructions for text evaluation. The criteria can be designed according to different tasks, so it is more flexible and can assess from multiple aspects. Meanwhile, to evaluate the level of LLMs’ hallucination, a variety of methods have been developed as well. This section explores these methods using direct assessment. These approaches include prompting the LLMs to generate responses or using a classifier to decide whether the input contains hallucinations.
4.1.1 LLMs as tool
In this section, we introduce the methods that leverage LLMs as tools to conduct hallucination evaluation.
Following the works of Sec. 3, SCALE (Lattimer et al., 2023) focuses on long-form dialogue and uses the LLM as an NLI model to generate scores. It breaks long source documents into chunks to measure the supportive level of the source text span.  Chen et al. (2023b) leverage prompting methods to evaluate summarization task. It designs chain-of-thoughts (CoT) and sentence-by-sentence prompts on few-shot and zero-shot scenarios. TrueTeacher (Gekhman et al., 2023) elaborates a data augmentation method, which uses LLM to label the generated summaries from CNN/DM, as well as multilingual data.
Expanding to a larger scope of tasks, GPTScore (Fu et al., 2023) and G-Eval (Liu et al., 2023) both provide multi-faceted evaluation frameworks including consistency. GPTScore (Fu et al., 2023) covers a range of tasks including dialogue, summarization, data-to-text, and NMT. When evaluating different tasks, normally, LLMs will be fed with the task introduction and the evaluation criteria as a prompt. G-Eval (Liu et al., 2023) uses CoT and a form-filling paradigm. When evaluating, it is asked to generate a CoT of evaluation steps formatted as a form. It shows that CoT can help GPT-4 better understand the evaluation criteria. Other than these two works, Wang et al. (2023b) study the feasibility of ChatGPT as a good evaluator for three tasks. This paper shows us a way to build evaluators based on ChatGPT with or without reference. ChainPoll (Friel and Sanyal, 2023) also uses ChatGPT to build a hallucination detection method with the CoT paradigm and considers it to be a binary classification problem.
Using LLMs for evaluation is simple and convenient. Despite the increasing trust in LLMs as they grow in size and capability, determining how to apply LLMs effectively to ensure their stability and reliability remains an open question.
4.1.2 LLMs as Object
In this section, we present recent evaluators and benchmarks for evaluating the hallucination level of LLMs from many significant aspects.
General Fact: TruthfulQA (Lin et al., 2022) is a benchmark to measure whether a language model is truthful enough. It brings up three concerns of the hallucination in the QA system: accidental misuse, blocking positive applications, and malicious misuse, and hedging is an acceptable solution for LLMs to be truthful in this work. HaluEval (Li et al., 2023) finds that ChatGPT is likely to generate unverifiable hallucinations and hard to identify the hallucinations in the generated text. Experiments suggest that extra knowledge and reasoning steps for LLMs can help alleviate hallucinations. FACTOR (Muhlgay et al., 2023) can measure LLM factuality concerning a controlled set of facts in a generation setting. It follows the error types from FRANK (Pagnoni et al., 2021) and performs a multi-choice factual evaluation task. Benefiting from the open source of recent LLMs, it is possible to access the parameters and dense distribution. EigenScore (Chen et al., 2024) exploits dense semantic information of LLMs to measure semantic consistency and diversity. Based on the observation of the existing extreme features in the LLM penultimate layer, it also use Test Time Feature Clipping to prevent LLMs generate overconfident hallucinations.
Furthermore, for evaluating the metrics themselves, RealHall (Friel and Sanyal, 2023) is a benchmark for evaluation methods and contains both closed- and open-domain hallucinations, which correspond to SF and WF. FELM (Zhao et al., 2024) also provides a meta-evaluation for metrics. It broadens the conventional understanding of world knowledge factuality to diverse domain data including world knowledge, science, math, recommendation, and reasoning. It uses ChatGPT to generate the responses for prompts and do segmentation, while humans are asked to label factuality and error reasons, error types, and references for each sample.
Fresh Fact: The world is constantly changing. How can we determine whether LLMs possess dynamic knowledge? The following benchmarks concentrate on constructing time-sensitive datasets to enable the evaluation of LLMs’ capacity to incorporate up-to-date information. FreshQA (Vu et al., 2023) is a QA benchmark that includes questions about current events and also inputs with false premises to the LLMs. The questions are divided into four types according to the freshness referring to the ever-changing world, including never-, slow-, fast-changing, and false-premise questions. RealTimeQA (Kasai et al., 2024) designs an annotation framework and submission workflow to support the updated evaluation system. It tests on both open- and closed-book QA systems and shows the importance of QA systems to be keep-dated. ERBench (Oh et al., 2024) verifies facts through entity-relationship (ER) model. It can construct questions using templates from updated ER databases and evaluates LLMs with binary or multi-choice QA tasks.
Long Context/Generation: Despite recent advancements in LLMs enabling them to better handle long texts, hallucination evaluation in long context or generation remains a challenge. FACTSCORE (Min et al., 2023) focuses on long-form text generated by LLMs. It breaks the generated text into a series of fine-grained atomic facts and then each atomic fact is assigned a binary label. The data is about biographies of people sampled from English Wikipedia. It refers to a given source of knowledge, which avoids information conflictions between different sources, but challenges the comprehensiveness and accuracy of the knowledge data. BAMBOO (Dong et al., 2023) provides a multi-task long context benchmark for LLMs which also can detect hallucination, which uses an entailment model on two synthetic datasets.
Domain Specific: Hallucinations generated by LLMs in the medical field can have serious consequences in healthcare applications. Therefore MedHalt (Pal et al., 2023) designs a pipeline to implement three fact-related tests: False Confidence Test (FCT), None of the Above Test (Nota), and Fake Questions Test (FQT). Through these tests, MedHalt can effectively identify and filter hallucinations within the medical domain.
Chinese Language: Numerous Chinese LLMs have also emerged along with the trend and hallucination is also a crucial problem. ChineseFactEval (Wang et al., 2023a) conducts the evaluation in seven scenarios, covering areas where generative AI may be applied in daily life, and specifically includes the modern Chinese history scenario. UHGEval (Liang et al., 2023) provides a Chinese-language hallucination dataset, generated by Chinese LLMs in news domain. The evaluation method uses LLMs for discriminative, selective and generative evaluation. HalluQA (Cheng et al., 2023) is a benchmark categorizing hallucinations into imitative falsehoods and factual errors. Inspired by TruthfulQA, it summarizes the question patterns and combines them with Chinese culture.
It should be noted that there is a small portion of overlap between Sec. 4.1.2 and Sec. 4.1.1, meaning that some approaches utilize LLMs to evaluate LLMs (Zhao et al., 2024) as well. For instance, due to its outstanding performance, GPT-4 can sometimes do annotation instead of human (Cheng et al., 2023). In contrast to evaluators before LLMs that primarily focus on SF, methods outlined in this section not only can address SF evaluation but also pay more attention to the WF (general fact) aspect, according to Figure 1. This is because LLMs enable open-domain QA and can even leverage search engines for retrieval. Thus, assessing fresh facts is also an important aspect. Though the latest LLMs can support long-text processing, challenges still persist in analyzing long contexts or generating long outputs, leading to hallucinations caused by long-distance discourse. How to catch the discourse-level errors while maintaining the ever-changing context is still challenging. Furthermore, in other application domains such as the medical field and non-English language, the collection of domain-specific data and the design of methodologies require further investigation, particularly in unique contexts such as medical statistics, legal terminology, and historical culture.
4.2 Pipeline Evaluation
Apart from utilizing LLMs for direct assessment, some evaluation methods do not rely on specific knowledge datasets or the pre-existing knowledge of LLMs. Instead, they leverage relevant evidence acquired through LLMs or search engines, benefitting from their enhanced search capabilities. Additionally, some methods involve the model generating answers multiple times to cross-check consistency to get the assessment outcome.
In a pipeline of fact extraction, evidence confirmation and fact consistency verification, FacTool (Chern et al., 2023) and UFO (Huang et al., 2024) both show competitive evaluation frameworks for hallucination. FacTool is a tool-augmented framework that includes includes multiple tasks and domains, which makes it relatively holistic. UFO provides a factuality verification method with multiple fact sources including human-written evidence, reference documents, search engine results, and LLM knowledge. CONNER (Chen et al., 2023a) also relies on external evidence to test LLMs the world knowledge they possess. It performs knowledge-intensitive tasks, including open-domain question answering and knowledge-grounded dialogue.
Utilizing LLM’s inherent capabilities for consistency checking is another viable approach. SelfCheckGPT (Manakul et al., 2023b) is motivated by the simple idea that LLM tends to have self-consistency if it possesses certain knowledge. It uses the scheme of black-box zero-resource hallucination detection and compares multiple sampled responses then measuring consistency. Rather than self-consistency, ${}\mathbf{SAC^{3}}$ (Zhang et al., 2023a) performs cross-checking hallucination detection based on sampling. By leveraging semantically equivalent question perturbation and cross-model response consistency checking, it can effectively identify the consistency and factuality of the answers. KoLA (Yu et al., 2023) use both Wikipedia and continuously collected emerging news and novels to evaluate the world knowledge of LLMs. It develops a self-contrast metric by contrasting two completions generated by the same model and gets the similarity score. RV (Yang et al., 2023) performs zero-resource hallucination detection by reverse validation. It focuses on passage-level detection rather than sentence-level, which holds greater practical significance. To evaluate the factual reasoning abilities of LLMs, Laban et al. (2023) build a benchmark SUMMEDITS. This work provides a three-step protocol for inconsistency detection benchmark creation and implements it in a 10-domain benchmark. It shows current LLMs are not yet proficient at complex factual reasoning.
These methods employ LLMs as tools to assess the hallucination level of models, including the LLMs themselves. While using multi-step approaches for evaluation offers strong interpretability, it is prone to error accumulation throughout extraction, reasoning, or multi-hop question answering. Similarly, having the model generate multiple answers or using multiple models to generate answers relies heavily on the models’ confidence and robustness. This dependency can lead to the oversight of high-confidence errors. We additionally present four case examples in Appendix D, covering both SF errors and WF errors. Table 2 shows that perfect evaluation for different types of hallucination errors remains challenging, and the evaluators also possess varying degrees of transferability.
5 Conclusion and Future Direction
Evaluating hallucination in NLG is crucial as it determines the direction and trend for developing better models. Historically, most metrics have been task-based, as defining the criteria for task performance is relatively straightforward. However, as NLP research gradually shifts to LLMs, researchers have proposed new demands and challenges. We summarize several aspects of hallucination evaluation that urgently need to be addressed in the future.
Hallucination and text error It can be challenging to distinguish between hallucinations and text errors. According to the traditional definition of hallucination (smooth but incorrect), any response from a large model that differs from the ground truth can be considered as hallucination, which is obviously unreasonable and can mislead researchers. Some works in NMT have already made progress in this area (Dale et al., 2023), and future evaluation methods should aim to accurately identify real hallucinations.
Hallucination and imagination Is hallucination always bad? Not necessarily. In certain contexts, such as chatting about a sci-fi novel, the content is rich with imagination, and the dialogue should be creative as well. In this circumstance, the distinction between hallucination and imagination is quite subtle. Distinguishing between these two phenomena can help models evaluate more diverse texts.
Supporting theories So far, the evaluation and detection of hallucination have been limited to observing the final output of the model – detecting hallucination manifested in the text. If there is theoretical support for researchers to detect hallucination from internal hidden states, model structures, and even data distributions, it can greatly reduce the cost of model iteration.
Interpretability Identifying fact granularity and analyzing the reasons behind hallucination can provide significant assistance in solving hallucination problems. Some reasoning methods have the potential to analyze the underlying causes of hallucinations and offer better evaluation. Other observing aspects lie in the internal state of model generation (Su et al., 2024), which provide more analytical perspectives.
Long context/generation Current exploration of LLMs in long-context scenarios mainly focuses on long inputs, such as using a complete contract or multi-turn dialogue history as input. It is crucial to address hallucinations caused by the model’s difficulty in understanding long inputs. However, hallucinations caused by contradictions between the beginning and end of long outputs are also worth exploring (Wei et al., 2024), such as detecting whether a model-generated novel contains inconsistencies in character behavior.
More application views LLMs have more challenging hallucinations in certain scenarios. In the multilingual scenario, low-resource languages other than English may be more prone to hallucinations, and cross-lingual conversation can face misunderstanding hallucinations due to different cultural backgrounds. Moreover, the latest research focuses on expanding LLMs to areas such as autonomous agents, multimodality, and real-world applications, which bring about new types of hallucinations, such as code hallucination (Qian et al., 2023), tool hallucination, visual hallucination, multimodal hallucination, and so on. Detecting such hallucinations is a very interesting and worthwhile direction to explore.
6 Limination
In this paper, we collect a broad range of related papers and reports, categorized and compared the methods and offered our insights for future directions. However, this paper does have some limitations. First of all, we did not do comprehensive experiments to revisit the above evaluators, because the evaluators usually focus on different types of hallucinations for various tasks, and it wouldn’t be fair to compare across the categories. For example, evaluators for LLMs intend to build their own datasets with human annotation, which vary in categories and schemes. Second, we put the content related to fact-checking and human evaluation in Appendix AB. We do not have enough pages to provide a thorough explanation of each of them, and detailed descriptions could also blur the focus of this paper. At last but not least, the case study we provide in Appendix D only includes a few representative cases on selective models for reference.
Acknowledgments
Appendix A Fact-checking
Fact-checking or fact-verification task is another line of work that has been paid much attention. The fact-checking framework can be divided into three components: claim detection, evidence retrieval, and claim verification (Guo et al., 2022), which is a relatively mature pipeline. Distinct from the evaluation methods discussed before in this paper, it typically involves assessing the factual accuracy of individual claims, mostly focusing on their WF. Wikipedia is a commonly used source for world knowledge (Thorne et al., 2018; Schuster et al., 2021; Kamoi et al., 2023; Gupta et al., 2022; Schuster et al., 2021), not only for fact-checking, but also for factuality evaluation. Especially when extracting evidence from a specific source, the WF turns into SF, which also demonstrates the dialectical unity of WF and SF. Benefiting from LLMs, fact-checking can process longer and more complex texts with more confidence and efficiency (Wang et al., 2023c). Due to the nature of the fact-checking task, it can be seen as a WF evaluator for text generation with a binary (true/false) checker.
Appendix B Human Evaluation
For hallucination evaluation, human perspectives can play a pivotal role, providing datasets and establishing benchmarks for the development of automatic models. To build a human annotation framework, there are three aspects requiring consideration: 1) How to design error categories and unify guidelines for annotators; 2) How to ensure the reliability of human annotation; And 3) how to digitally present annotated results. Human evaluation can be time-consuming and is particularly inefficient for large-scale evaluations, but still is the most trustworthy way of model evaluation.
Appendix C Supplement Table
We provide a table of evaluators here, as in Table 1. For the Based-model column, it means the models that evaluators use to perform evaluation or generate synthetic data. Metric column means the calculating method to get the final score. ✓and ✗ in SF and WF columns mean the aspects that the evaluators focus on.
Appendix D Case Study
Here we present some results of selected evaluators on different kinds for SF or WF errors on summarization data in Table 2. The data we used are from XEnt dataset (Cao et al., 2022) and FactCollect (Ribeiro et al., 2022). We selected evaluators that use the GPT series and those that do not, covering both models that evaluate SF and WF facets. For the models utilizing LLMs, we specifically employed GPT-3.5-turbo. Although FacTool is not directly applicable for evaluating summarization tasks, we conducted experiments under its KBQA (Knowledge-Based Question Answering) setup to see its transfer ability.
The results of different models on these cases show considerable variation. In the SF-WF case, only FacTool made an incorrect judgment, which might be attributed to its insufficient transfer ability. SelfCheckGPT uses a zero-shot approach in its prompt to assess the consistency, whereas HaluEval’s prompt provides examples for judgment. However, the SFE cases indicate that the results of these two evaluators remain unstable. For the WFE cases, FacTool provides the correct answers, and surprisingly, WeCheck also made correct judgments. Currently, to our best knowledge, no such labeled data is available for full evaluation. More accurate data is needed for further experiments to validate the preferences of different evaluators.
Era Name New Dataset Data Source Fact Definition Task Based-model Method Metric SF WF Before LLM Ref -based $Fact_{acc}$ WikiFact Wikipedia, Wikidata KB Triplet Summ Transformer NER, Classification P, R, F1 FactCC FactCC CNN/DM, XSumFaith Sent Summ BERT NLI (2-class) Likelihood DAE DAE PARANMT50M Dependency Summ ELECTRA NLI (2-class) Likelihood Maskeval / CNN/DM, WikiLarge, ASSET Word Summ, Simp T5 Word Weighting Weighted Match Score  Guerreiro et al. (2023) Haystack WMT2018, DE-EN Text Span NMT Transformer Uncertainty Measure Avg. Similarity  Dale et al. (2023) / Haystack Text Span NMT Transformer Source Contribution Percentage Before LLM Ref -free FEQA FEQA CNN/DM, XSum Sent Span Summ BART (QG), BERT (QA) QG-QA Avg. F1 QAGS QAGS CNN/DM, XSum Ent, Noun Phrase Summ BART (QG), BERT (QA) QG-QA Avg. Similarity QuestEval / CNN/DM, Xsum Ent, Noun Summ T5 (QG, QA) QG-QA P, R, F1 QAFactEval / SummaC NP Chunk Summ BART (QG), ELECTRA (QA) QG-QA, NLI LERC MQAG / QAGS, XSumFaith, Podcast, Assessment, SummEval Sent Span Summ T5 (QG), Longformer (QA) Multi-Choice QA Choice Statistical Distance CoCo / QAGS, SummEval Token, Span, Sent, Doc Summ BART Counterfactual Estimation Avg. Likelihood Diff FactGraph FactCollect CNN/DM, XSum Dependency Summ ELECTRA Classification BACC, F1 FactKB FactKB CNN/DM, XSum Triplet Summ RoBERTa Classification BACC, F1 ExtEval ExtEval CNN/DM Discourse, Coreference, Sentiment Summ SpanBERT, RoBERTa Direct Prediction, Statistic Summation of Sub-scores $Q^{2}$ $Q^{2}$ WOW Sent Span Diag T5 (QG), Albert-Xlarge (QA), RoBERTa (NLI) QG-QA, NLI Likelihood FactPush / TRUE Span Diag, Summ, Paraphrase DeBERTa NLI AUC AlignScore / 22 datasets from 7 tasks Sent NLI, QA, Paraphrase, Fact Verification, IR, Semantic Similarity, Summ RobERTa 3-way Classification Likelihood WeCheck / TRUE Response Summ, Diag, Para, Fact Check DeBERTaV3 Weakly Supervised NLI Likelihood After LLM Direct SCALE ScreenEval LLM, Human Sentence Long Diag Flan-T5 NLI Likelihood  Chen et al. (2023b) / SummEval, XSumFaith, Goyal21, CLIFF Response Summ Flan-T5, code-davinci-002, text-davinci-003, ChatGPT, GPT-4 Vanilla/COT/ Sent-by-Sent Prompt Balanced Acc GPTScore / 37 datasets from 4 tasks Various Summ, Diag, NMT, D2T GPT-2, OPT, FLAN, GPT-3 Direct Assessment Direct Score G-Eval / SummEval, Topical-Chat, QAGS Response Summ, Diag GPT-4 COT, Form-filling Weighted Scores Wang et al. (2023b) / 5 datasets from 3 tasks Response Summ, D2T, Story Gen ChatGPT Direct Assessment, Rating Direct score ChainPoll RealHall-closed, RealHall-open COVID-QA, DROP, Open Ass prompts, TriviaQA Response Hallu Detect gpt-3.5-turbo Direct Assessment (2-class) Acc EigenScore / CoQA, SQuAD, TriviaQA Natural Questions Inner State Open-book QA Closed-book QA LLaMA, OPT Semantic Consistency/ Diversity in Dense Embedding Space AUROC, PCC TruthfulQA TruthfulQA LLM, Human Response Multi-Choice QA, Generation GPT-3-175B Answer Match Score HaluEval Task-specific, General Alpaca, Task datasets ChatGPT Response QA, Summ, Knowledge- grounded Diag, Generation ChatGPT Direct Assessment Acc FACTOR Wiki-/News-/ Expert- FACTOR Wikipedia, Refin- edWeb, ExpertQA Sent Span Generation / FRANK Error Classification likelihood FELM FELM TruthfulQA, Quora, MMLU, GSM8K, ChatGPT, Human Text Span, Claim World Knowledge, Sci and Tech, Math, Writing and Recom- mendation, Reasoning Vicuna, ChatGPT, GPT4 Direct Assessment F1, Balanced Acc FreshQA Never/Slow Fast-changing, false-premise Human Response Generation / Answer Match Acc RealTimeQA RealTimeQA CNN, THE WEEK, USA Today Response Multi-Choice QA, Generation GPT-3, T5 Answer Match Acc, EM, F1 ERBench ERBench Database 5 datasets from Kaggle Ent-Rel Binary/ Multiple -choice QA / Direct Assessment, String Matching Ans/Rat/ Ans-Rat Acc, Hallu Rate FactScore / Biographies in Wikipedia Atomic Fact Generation InstructGPT, ChatGPT, PerplexityAI Binary Classification P BAMBOO SenHallu, AbsHallu 10 datasets from 5 tasks Response Multi-choice tasks, Select tasks ChatGPT Answer Match P, R, F1 MedHalt MedHalt MedMCQA, Medqa USMILE, Medqa (Taiwan), Headqa, PubMed Response Reasoning Hallu Test, Memory Hallu Test ChatGPT Answer Match Pointwise Score, Acc ChineseFactEval ChineseFactEval / Response Generation / FacTool, Human annotator Direct Score UHGEval UHGEval Chinese News Websites Keywords Generative/ Discriminative/ Selective Evaluator GPT-4 Answer Match, Similarity Acc, Similarity Score HalluQA HalluQA Human Response Generation GLM-130B, ChatGPT, GPT-4 Direct Assessment Non-hallu Rate After LLM Pipeline FacTool / RoSE, FactPrompts, HumanEval, GSM-Hard, Self-instruct Claim, Response Knowledge-based QA, Code Generation, Math Reasoning, Sci-literature Review ChatGPT Claim Extraction, Query Generation, Tool Querying, Evidence Collection, Agreement Verification P, R, F1 UFO / NQ, HotpotQA, TruthfulQA, CNN/DM, Multi-News, MS MARCO Ent Open-domain/ Web Retrieval-based/ Expert-validated/ Retrieval-Augmented QA, News Fact Generation ChatGPT (gpt-3.5-turbo-1106) Fact Unit Extraction, Fact Source Verification, Fact Consistency Discrimination Avg. Sub-scores CONNER / NQ, WoW Sentence Open-domain QA, Knowledge-grounded Dialogue NLI-RoBERTa -large, ColBERTv2 3-way NLI Acc SelfCheckGPT SelfCheckGPT WikiBio Response Hallu Detect GPT-3 NLI, Ngram, QA, BERTScore, Prompt AUC-PR $SAC^{3}$ / HotpotQA, NQ-open Response QA Generation gpt-3.5-turbo, Falcon-7b-instruct, Guanaco-33b Cross-checking, QA Pair Consistency AUROC KoLA KoLA Wikipedia, Updated News and Novels Response Knowledge Memorization /Understanding/Applying /Creating / Self-contrast Answer Match Similarity RV PHD Human Annotator Ent Generation ChatGPT Construct Query, Access Databases, Entity-Answer Match P, R, F1 SummEdits SummEdits 9 datasets from Summ task Span Summ, Reasoning gpt-3.5-turbo Seed summary verify, Summary edits, Annotation Balanced Acc
Table 1: Evaluators and Benchmarks meta information.
Document Summary Note WeCheck SelfCheckGPT HaluEval FacTool SF-WF … Harry Kane has been given the nod by Youssouf Mulumbu for this season’s players’ Player of the Year award The West Brom midfielder has picked Chelsea wideman Eden Hazard for the young player of the year prize Congo international Mulumbu posted his votes for this year’s PFA awards to Twitter on Wednesday Mulumbu challenges QPR defender Yun Suk-Young during West Brom’s 4-1 defeat at The Hawthorns Goalkeepe … The DR Congo international has picked Chelsea wideman Eden Hazard for the young player of the year prize . The summary is correct. TRUE TRUE TRUE FALSE SF-WFE … Since the end of March, the Vikings’ only wins have been in the Challenge Cup against lower-league sides. "We’ve got the personnel and we’ve got the people to spark us back into life," Chris Betts told BBC Radio Merseyside. "When we get rolling again I’m sure, or I’m positive, that we can really turn this year around for ourselves." … "The players are hurting and we’ve got to win," added England assistant coach Betts. … Widnes Vikings can turn their poor start to the Super League season around if they can find a winning streak, says assistant coach Chris Betts. "Chris Betts" is in the document but is incorrect essentially. FALSE TRUE TRUE FALSE SFE-WF The panther chameleon was found on Monday by a dog walker in the wooded area at Marl Park. It had to be put down after X-rays showed all of its legs were broken and it had a deformed spine. RSPCA Cymru said it was an "extremely sad example of an abandoned and neglected exotic pet". …… A chameleon has been put down by RSPCA Cymru after it was found injured and abandoned in a Cardiff park. The Marl Park is in Cardiff but not mentioned in the document. TRUE FALSE TRUE TRUE SFE-WFE A number of men, two of them believed to have been carrying guns, forced their way into the property at Oakfield Drive shortly after 20:00 GMT on Saturday. They demanded money before assaulting a man aged in his 50s. … Alliance East Antrim MLA Stewart Dickson has condemned the attack. … A man has been assaulted by a gang of armed men during a robbery at a house in Ballymena, County Antrim. "Ballymena" is neither in the document nor correct according to external knowledge. FALSE TRUE TRUE FALSE
Table 2: Examples of the results from selected evaluators on the SF errors (SFE) and WF errors (WFE).